OpenStudy (anonymous):

The given equation is either linear or equivalent to a linear equation. Solve the equation. u/(u-(u+1)/2)=4

OpenStudy (anonymous):

\[\frac{u}{\frac{u-(u+1)}{2}}=4\]?

OpenStudy (anonymous):

looks like you have two fraction bars there.

OpenStudy (anonymous):

yes satellite. haha

myininaya (myininaya):

\[\frac{u}{u-\frac{(u+1)}{2}}=4\]

OpenStudy (anonymous):

ok well first of all \[u-(u+1)=u-u-1=-1\]

OpenStudy (anonymous):

so what you really have is \[\frac{u}{\frac{-1}{2}}=4\]

OpenStudy (anonymous):

dividing by one half same as multiplying by 2 so you get \[-2u=4\]

OpenStudy (anonymous):

divide by -2 to get \[u=\frac{4}{-2}=-2\]

OpenStudy (anonymous):

no. stop. sorry. wait a sec

myininaya (myininaya):

lol

OpenStudy (anonymous):

i stop

OpenStudy (anonymous):

myininya has the correct problem shown

myininaya (myininaya):

:)

OpenStudy (anonymous):

well then my answer was the answer to a different question!

OpenStudy (anonymous):

;(

OpenStudy (anonymous):

using this site as a reference: http://www.wolframalpha.com/input/?i=u%2F%28u-%28u%2B1%29%2F2%29%3D4

OpenStudy (anonymous):

any input is appreciated. i'm trying to figure out how to work it...

OpenStudy (anonymous):

well that certainly is not a linear equation!

myininaya (myininaya):

i would combine the fractions on bottom first

OpenStudy (anonymous):

The given equation is either linear or equivalent to a linear equation. Solve the equation. u/(u-(u+1)/2)=4

OpenStudy (anonymous):

no chance. it is not linear

OpenStudy (anonymous):

haha. that's how it's listed in my book :D

OpenStudy (anonymous):

well it isn't look at the graph!

OpenStudy (anonymous):

are you sure you typed it in correctly?

OpenStudy (anonymous):

u = 2

OpenStudy (anonymous):

yes. look at the link. that's exactly as it appears in my book :)

OpenStudy (anonymous):

sure it isn't -2?

OpenStudy (anonymous):

it's 2

OpenStudy (anonymous):

aaaaaaaaaaaaaaaaaahhhhhhhhhhhhhhhhhhhhhhh

OpenStudy (anonymous):

i see. simplify the compound fraction and you will get 2 lickety split

myininaya (myininaya):

\[\frac{u}{\frac{2u-(u+1)}{2}}=u \div \frac{(2u-(u+1))}{2}=u \times \frac{2}{u-1}=\frac{2u}{u-1}\]

OpenStudy (anonymous):

i don't see... :(

myininaya (myininaya):

thats the left hand side

myininaya (myininaya):

\[\frac{2u}{u-1}=4\]

OpenStudy (anonymous):

where did you get 2u from?

OpenStudy (anonymous):

in the very first step...

OpenStudy (anonymous):

ok multiply top and bottom by 2 get \[\frac{2u}{2u-(u+1)}\]

OpenStudy (anonymous):

no pellet... lol. okay. i'm with you

OpenStudy (anonymous):

clears the annoying 2 in the bottom of the denominator

OpenStudy (anonymous):

now as we say in the math biz combine liked terms to get \[\frac{2u}{u-1}\]

myininaya (myininaya):

satellite cheated and did it the easy way

OpenStudy (anonymous):

yeah. he thinks i know things and understand. :p

myininaya (myininaya):

lol

OpenStudy (anonymous):

so we have \[\frac{2u}{u-1}=4\] \[2u=4(u-1)\] \[2u=4u-4\] \[u = 2\] whew

OpenStudy (anonymous):

waddya mean cheated? how else do you clear fractions?

OpenStudy (anonymous):

wow. EASY!!! why was i stressing that?

myininaya (myininaya):

lol i was kidding satellite

OpenStudy (anonymous):

because it is math

OpenStudy (anonymous):

i was taught to multiply the whole left side by the complex fraction u-(u+1/2) and do the same on the right...

myininaya (myininaya):

math is the best because it is logic english makes no sense

OpenStudy (anonymous):

haha. taking math and english.... smh...

OpenStudy (anonymous):

ooohhhh i see you subtracted in the denominator then inverted and multiplied. ok that works too.

OpenStudy (anonymous):

if you are into long algebra. already annoyed that you can out-latex me! lol

OpenStudy (anonymous):

gonna grab some food. will be back to annoy you guys in a little ;) thank you both a lot!

OpenStudy (anonymous):

now you are going to out algebra me too. not fair

myininaya (myininaya):

lol no shorter ways are better than longer ways lol i can only out latex you when you are drunk